# Learning nonnegative matrix factorizations from compressed data

### v1

@@ 3.1. Two-sided compression
$P_{{\mathbf{A}}_{2}^{T}{\mathbf{X}}^{T}}^{\perp}\tilde{\mathbf{U}}\tilde{\mathbf{V}}^{T}=\tilde{\mathbf{U}}\tilde{\mathbf{V}}^{T}P_{{\mathbf{A}}_{1}{\mathbf{X}}}^{\perp}=\mathbf{0}.$ (3.3)
By (3.3), we can write
$\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}=\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}P_{{\mathbf{A}}_{1}{\mathbf{X}}}=P_{{\mathbf{A}}_{2}^{T}{\mathbf{X}}^{T}}\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}P_{{\mathbf{A}}_{1}{\mathbf{X}}}={\mathbf{X}}{\mathbf{A}}_{2}{\mathbf{M}}{\mathbf{A}}_{1}{\mathbf{X}},$ (3.4)
@@
${\mathbf{M}}:=(({\mathbf{X}}{\mathbf{A}}_{2})^{T}{\mathbf{X}}{\mathbf{A}}_{2})^{-1}({\mathbf{X}}{\mathbf{A}}_{2})^{T}\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}({\mathbf{A}}_{1}{\mathbf{X}})^{T}({\mathbf{A}}_{1}{\mathbf{X}}({\mathbf{A}}_{1}{\mathbf{X}})^{T})^{-1}.$
Now,
@@
and since ${\mathbf{A}}_{1}{\mathbf{X}}{\mathbf{A}}_{2}$ is invertible we have ${\mathbf{M}}=({\mathbf{A}}_{1}{\mathbf{X}}{\mathbf{A}}_{2})^{-1}$. Thus, (3.4) implies
$\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}={\mathbf{X}}{\mathbf{A}}_{2}({\mathbf{A}}_{1}{\mathbf{X}}{\mathbf{A}}_{2})^{-1}{\mathbf{A}}_{1}{\mathbf{X}}={\mathbf{Y}}={\mathbf{X}}.$
@@ 3.2. One-sided compression: orthogonal sketching matrices
Theorem 3.4.
(Orthogonal ${\mathbf{A}}$) Let ${\mathbf{X}}\in{\mathbb{R}}_{+}^{m\times n}$ be any matrix and let ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ be a matrix with orthogonal rows (i.e. ${\mathbf{A}}{\mathbf{A}}^{T}={\mathbf{I}}$). Let ${\mathbf{U}}_{0}\in{\mathbb{R}}_{+}^{m\times r}$, ${\mathbf{V}}_{0}\in{\mathbb{R}}_{+}^{r\times n}$ give a solution to the original NMF problem (1.1) of rank $r$ and ${\mathbf{X}}_{0}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}$. If $\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}$ solve a compressed NMF problem with the same rank $r$, that is,
$\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}=\operatorname*{arg\,min}_{{\mathbf{U}},{\mathbf{V}}\geq 0}\left[\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|^{2}_{F}\right],$ (3.5)
where $\lambda>0$, $P_{\mathbf{A}}^{\perp}:={\mathbf{I}}-{\mathbf{A}}^{T}{\mathbf{A}}$, and $({\mathbf{U}},{\mathbf{V}}\geq 0)$ means $({\mathbf{U}}\in{\mathbb{R}}_{+}^{m\times r},{\mathbf{V}}\in{\mathbb{R}}_{+}^{r\times n})$. Then $\tilde{{\mathbf{X}}}:=\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}$ satisfies
$\frac{\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}\leq c_{\lambda}\left[\frac{\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}+\frac{\|P^{\perp}_{\mathbf{A}}{\mathbf{X}}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}\right],$ (3.6)
@@
We note that this and further results do not require the data matrix to have an exact nonnegative low rank $r$ as in Theorem 3.1 (although if it is, the first term in the bound for the loss vanishes). Before we start the proof, let us recall a simple corollary of the Pythagorean theorem and the triangle inequality to be used several times below: for any matrices ${\mathbf{X}},{\mathbf{Y}}$ and a projection operator $P_{\mathbf{A}}$,
$\|{\mathbf{X}}-{\mathbf{Y}}\|^{2}_{F}$ $\leq\|P_{\mathbf{A}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}+2\|P_{\mathbf{A}}^{\perp}{\mathbf{Y}}\|^{2}_{F}+2\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}.$ (3.7)
@@
$\|{\mathbf{X}}-{\mathbf{Y}}\|^{2}_{F}$ $=\|P_{\mathbf{A}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}+\|P_{\mathbf{A}}^{\perp}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}$
$\leq\|P_{\mathbf{A}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}+(\|P_{\mathbf{A}}^{\perp}{\mathbf{Y}}\|_{F}+\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|_{F})^{2}$
$\leq\|P_{\mathbf{A}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}+2\|P_{\mathbf{A}}^{\perp}{\mathbf{Y}}\|^{2}_{F}+2\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}.$
Proof of Theorem 3.4.
@@
$\|{\mathbf{A}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}\leq\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{X}}_{0}^{T})\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}_{0}\|^{2}_{F},$ (3.8)
since $\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}$ minimize the objective of (3.5) over all nonnegative matrices of the appropriate sizes. Now, since ${\mathbf{A}}{\mathbf{A}}^{T}={\mathbf{I}}$, observe that for any ${\mathbf{M}}\in{\mathbb{R}}^{m\times n}$ matrix
$\|{\mathbf{A}}{\mathbf{M}}\|_{F}=\|{\mathbf{A}}^{T}{\mathbf{A}}{\mathbf{M}}\|_{F}=\|P_{\mathbf{A}}{\mathbf{M}}\|_{F}.$
So, using identity (3.7) for the matrices ${\mathbf{X}}$ and $\tilde{{\mathbf{X}}}$, we can estimate
$\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|^{2}_{F}$ $\leq\|{\mathbf{A}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+2\|P_{\mathbf{A}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}+2\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq c_{1}\big{(}\|{\mathbf{A}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}\big{)}+2\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
(3.8)${\leq}c_{1}\big{(}\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}_{0}\|^{2}_{F}\big{)}+2\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
for $c_{1}=\max(2/\lambda,1).$ Using identity (3.7) for the matrices ${\mathbf{X}}$ and ${\mathbf{X}}_{0}$, we can estimate the term in parentheses as
$\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}_{0}\|^{2}_{F}$ $\leq\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+2\lambda\|P_{\mathbf{A}}^{\perp}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+2\lambda\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq c_{2}\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}+2\lambda\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
@@
$\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|^{2}_{F}\leq c_{1}c_{2}\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}+(2\lambda c_{1}+2)\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}.$
With $c_{\lambda}=\max(c_{1}c_{2},2\lambda c_{1}+2)=\max(2/\lambda,6,2\lambda+2)$, this concludes the proof of Theorem 3.4. ∎
Theorem 3.4 shows that the solution to the compressed NMF problem (1.1) will work for the original uncompressed problem (3.5) as long as the terms $\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|^{2}_{F}$ and $\|P^{\perp}_{\mathbf{A}}{\mathbf{X}}\|^{2}_{F}$ are small. Luckily, with just one more pass over the original data one can get such sketching matrices using the standard approaches of randomized linear algebra, such as those in [14].
Theorem 3.5 (“Randomized rangefinder algorithm loss”, [14]).
@@
$\mathbb{E}\|{\mathbf{X}}-P_{{\mathbf{S}}^{T}{\mathbf{X}}^{T}}{\mathbf{X}}\|_{F}\leq\left(1+\frac{r}{k-r-1}\right)^{\frac{1}{2}}\left(\sum_{j>r}\sigma_{j}^{2}({\mathbf{X}})\right)^{\frac{1}{2}},$
@@
Corollary 3.6 (Data-adapted one-sided sketches).
@@
Take $k$ such that $2r+1\leq k\leq\min\{m,n\}$. Form a sketch ${\mathbf{X}}{\mathbf{S}}$ with ${\mathbf{S}}$ is an $n\times k$ standard Gaussian matrix; find ${\mathbf{Q}}$, orthonormal basis of the column space of ${\mathbf{X}}{\mathbf{S}}$; take a sketching matrix ${\mathbf{A}}:={\mathbf{Q}}^{T}$. If $\tilde{{\mathbf{U}}}\in{\mathbb{R}}_{+}^{m\times r},\tilde{{\mathbf{V}}}\in{\mathbb{R}}_{+}^{r\times n}$ solve a compressed NMF problem (3.5) with this ${\mathbf{A}}$ and some $\lambda>0$, then
$\frac{{\mathbb{E}}\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|_{F}}{\|{\mathbf{X}}\|_{F}}\leq\sqrt{2}c_{\lambda}\varepsilon$ (3.9)
and $c_{\lambda}$ is the constant from (3.6).
@@
By Theorem 3.5 and approximate low-rankness of ${\mathbf{X}}$, we have
$\mathbb{E}\|{\mathbf{X}}-P_{\mathbf{A}}{\mathbf{X}}\|_{F}\leq\sqrt{1+\frac{r}{k-r-1}}\sqrt{\sum_{j>r}\sigma_{j}({\mathbf{X}})}\leq\sqrt{2}\|{\mathbf{X}}-{\mathbf{X}}_{0}\|_{F}\leq\sqrt{2}\varepsilon\|{\mathbf{X}}\|_{F},$
using that $k\geq 2r+1$ in the second inequality. Combining this with Theorem 3.4, we have
${\mathbb{E}}\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|_{F}\leq c(\|{\mathbf{X}}-{\mathbf{X}}_{0}\|_{F}+{\mathbb{E}}\|{\mathbf{X}}-P_{\mathbf{A}}{\mathbf{X}}\|_{F})\leq c(1+\sqrt{2})\varepsilon\|{\mathbf{X}}\|_{F},$
where $c=\max(2/\lambda,6,2\lambda+2).$
Remark 3.7.
A high probability deviation bound for the loss $\|{\mathbf{X}}-{\mathbf{Q}}{\mathbf{Q}}^{T}{\mathbf{X}}\|_{F}$ is also known [14, Theorem 10.7]. It implies a high probability estimate for (3.9) in a straightforward way. Instead of Gaussian initial sketching, one can employ subsampled random Fourier transform [14] or other cost-efficient matrices [31, 10].
It is easy to see that the oversampling condition $k>2r$ can be relaxed to any $k>r+1$ by suitably increasing the constant factor $\sqrt{2}$. Notwithstanding this factor, we see that if ${\mathbf{X}}$ has an exact NMF decomposition of rank $r$ and $k>r+1$ then the error of the optimal solution to the compressed problem must be also zero, comparable with the result of Theorem 3.1.

### v2

@@ 3.1. Two-sided compression
(4) $P_{{\mathbf{X}}{\mathbf{A}}_{2}}^{\perp}\tilde{\mathbf{U}}\tilde{\mathbf{V}}^{T}=\tilde{\mathbf{U}}\tilde{\mathbf{V}}^{T}P_{{\mathbf{X}}^{T}{\mathbf{A}}_{1}^{T}}^{\perp}=\mathbf{0}.$
Let ${\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}\Sigma{\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}^{T}$ be a compact singular value decomposition of $\mathbf{X}$} (i.e., such that $\Sigma$ is invertible). Observe that $P_{{\mathbf{X}}}={\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}{\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}^{T}$ and $P_{{\mathbf{X}}^{T}}={\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}{\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}^{T}$ and that the matrices ${\mathbf{A}}_{1}{\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}$ and ${\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}^{T}{\mathbf{A}}_{2}$ have the same rank as $\mathbf{X}$}. Then we can write
(5) $\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}=\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}P_{{\mathbf{X}}^{T}}=P_{{\mathbf{X}}}\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}P_{{\mathbf{X}}^{T}}={\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}{\mathbf{M}}{\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}^{T},$
@@
${\mathbf{M}}:={\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}^{T}\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}{\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}.$
Now observe
@@
The rank of ${\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}^{T}{\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}{\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}$ is the same as the rank of ${\mathbf{A}}_{1}{\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}$ which is the same as the rank of $\mathbf{X}$}, which is the same as the dimensions of ${\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}^{T}{\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}{\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}$ (since the SVD is compact); thus, the matrix ${\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}^{T}{\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}{\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}$ is invertible and likewise, the matrix ${\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}^{T}{\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}{\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}$ is invertible. Thus, we have
${\mathbf{M}}$ $=({\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}^{T}{\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}{\color[rgb]{0,0,0}\bar{{\mathbf{U}}}})^{-1}{\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}^{T}{\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}{\mathbf{X}}{\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}{\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}({\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}^{T}{\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}{\color[rgb]{0,0,0}\bar{{\mathbf{V}}}})^{-1}$
$=({\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}^{T}{\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}{\color[rgb]{0,0,0}\bar{{\mathbf{U}}}})^{-1}{\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}^{T}{\mathbf{A}}_{1}^{T}{\mathbf{A}}_{1}{\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}\Sigma{\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}^{T}{\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}{\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}({\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}^{T}{\mathbf{A}}_{2}{\mathbf{A}}_{2}^{T}{\color[rgb]{0,0,0}\bar{{\mathbf{V}}}})^{-1}$
$=\Sigma,$
and therefore we conclude that $\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}={\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}{\mathbf{M}}{\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}^{T}={\color[rgb]{0,0,0}\bar{{\mathbf{U}}}}\Sigma{\color[rgb]{0,0,0}\bar{{\mathbf{V}}}}^{T}={\mathbf{X}}$.
Remark 2.
Note that the conclusion of 1 holds for any matrices with zero objective value for problem (2), regardless of the constraint set — so, in particular, nonnegativity of the factors ${\mathbf{U}}_{0}$ and ${\mathbf{V}}_{0}$ is not used directly in the proof beyond the fact that the objective value of the uncompressed problem is zero.
@@ 3.2. One-sided compression: orthogonal sketching matrices
Theorem 2.
(Orthogonal $\mathbf{A}$}) Let ${\mathbf{X}}\in{\mathbb{R}}_{+}^{m\times n}$ be any matrix and let ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ be a matrix with orthogonal rows (i.e., ${\mathbf{A}}{\mathbf{A}}^{T}={\mathbf{I}}$). Let ${\mathbf{U}}_{0}\in{\mathbb{R}}_{+}^{m\times r}$, ${\mathbf{V}}_{0}\in{\mathbb{R}}_{+}^{r\times n}$ give a solution to the original NMF problem (1) of rank $r$ and ${\mathbf{X}}_{0}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}$. Suppose $\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}$ solve a compressed NMF problem with the same rank $r$, that is,
(6) $\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}{\color[rgb]{0,0,0}\;\in\;}\operatorname*{arg\,min}_{{\mathbf{U}},{\mathbf{V}}\geq 0}\left[\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2}_{F}+\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|^{2}_{F}\right],$
where $\lambda>0$, $P_{{\mathbf{A}}^{T}}^{\perp}:={\mathbf{I}}-{\mathbf{A}}^{T}{\mathbf{A}}$, and $({\mathbf{U}},{\mathbf{V}}\geq 0)$ means $({\mathbf{U}}\in{\mathbb{R}}_{+}^{m\times r},{\mathbf{V}}\in{\mathbb{R}}_{+}^{r\times n})$. Then $\tilde{{\mathbf{X}}}:=\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}$ satisfies
(7) $\frac{\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}\leq c_{\lambda}\left[\frac{\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}+\frac{\|P^{\perp}_{{\mathbf{A}}^{T}}{\mathbf{X}}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}\right],$
@@
Before we start the proof, let us recall a simple corollary of the Pythagorean theorem and the triangle inequality to be used several times below: for any matrices ${\mathbf{X}},{\mathbf{Y}}$ and a projection operator $P_{{\mathbf{A}}^{T}}$,
(8) $\|{\mathbf{X}}-{\mathbf{Y}}\|^{2}_{F}$ $\leq\|P_{{\mathbf{A}}^{T}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}+2\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{Y}}\|^{2}_{F}+2\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}.$
@@
$\|{\mathbf{X}}-{\mathbf{Y}}\|^{2}_{F}$ $=\|P_{{\mathbf{A}}^{T}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}+\|P_{{\mathbf{A}}^{T}}^{\perp}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}$
$\leq\|P_{{\mathbf{A}}^{T}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}+(\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{Y}}\|_{F}+\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|_{F})^{2}$
$\leq\|P_{{\mathbf{A}}^{T}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}+2\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{Y}}\|^{2}_{F}+2\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}.$
Proof of Theorem 2.
@@
(9) $\|{\mathbf{A}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}\leq\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{X}}_{0}^{T})\|^{2}_{F}+\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}_{0}\|^{2}_{F},$
since $\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}$ minimize the objective of (6) over all nonnegative matrices of the appropriate sizes. Now, since ${\mathbf{A}}{\mathbf{A}}^{T}={\mathbf{I}}$, observe that for any ${\mathbf{M}}\in{\mathbb{R}}^{m\times n}$ matrix
$\|{\mathbf{A}}{\mathbf{M}}\|_{F}=\|{\mathbf{A}}^{T}{\mathbf{A}}{\mathbf{M}}\|_{F}=\|P_{{\mathbf{A}}^{T}}{\mathbf{M}}\|_{F}.$
So, using inequality (8) for the matrices $\mathbf{X}$} and $\tilde{{\mathbf{X}}}$, we can estimate
$\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|^{2}_{F}$ $\leq\|{\mathbf{A}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+2\|P_{{\mathbf{A}}^{T}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}+2\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq c_{1}\big(\|{\mathbf{A}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}\big)+2\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
[…]
for $c_{1}=\max(2/\lambda,1).$ Using inequality (8) for the matrices $\mathbf{X}$} and ${\mathbf{X}}_{0}$, we can estimate the term in parentheses as
$\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}$ $+\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}_{0}\|^{2}_{F}$
$\leq\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+2\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+2\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq c_{2}\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}+2\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
@@
$\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|^{2}_{F}\leq c_{1}c_{2}\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}+(2\lambda c_{1}+2)\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}.$
With $c_{\lambda}=\max(c_{1}c_{2},2\lambda c_{1}+2)=\max(2/\lambda,6,2\lambda+2)$, this concludes the proof of Theorem 2. ∎
Theorem 2 shows that the solution to the compressed NMF problem (1) will work for the original uncompressed problem (6) as long as the terms $\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|^{2}_{F}$ and $\|P^{\perp}_{{\mathbf{A}}^{T}}{\mathbf{X}}\|^{2}_{F}$ are small. Luckily, with just one more pass over the original data one can get such sketching matrices using the standard approaches of randomized linear algebra, such as those in [15].
Theorem 3 (“Randomized rangefinder algorithm loss”, [15]).
@@
$\mathbb{E}\|{\mathbf{X}}-P_{{\mathbf{X}}{\mathbf{S}}}{\mathbf{X}}\|_{F}\leq\left(1+\frac{r}{k-r-1}\right)^{\frac{1}{2}}\left(\sum_{j>r}\sigma_{j}^{2}({\mathbf{X}})\right)^{\frac{1}{2}},$
@@
Corollary 1 (Data-adapted one-sided sketches).
@@
Take $k$ such that $2r+1\leq k\leq\min\{m,n\}$. Form a sketch ${\mathbf{X}}{\mathbf{S}}$ with $\mathbf{S}$} is an $n\times k$ standard Gaussian matrix; find $\mathbf{Q}$}, an orthonormal basis of the column space of ${\mathbf{X}}{\mathbf{S}}$; take a sketching matrix ${\mathbf{A}}:={\mathbf{Q}}^{T}$. If $\tilde{{\mathbf{U}}}\in{\mathbb{R}}_{+}^{m\times r},\tilde{{\mathbf{V}}}\in{\mathbb{R}}_{+}^{r\times n}$ solve a compressed NMF problem (6) with this $\mathbf{A}$} and some $\lambda>0$, then
(10) $\frac{{\mathbb{E}}\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|_{F}}{\|{\mathbf{X}}\|_{F}}\leq\sqrt{2}c_{\lambda}\varepsilon$
and $c_{\lambda}$ is the constant from (7).
@@
By Theorem 3 and approximate low-rankness of $\mathbf{X}$}, we have
$\mathbb{E}\|{\mathbf{X}}-P_{{\mathbf{A}}^{T}}{\mathbf{X}}\|_{F}\leq\sqrt{1+\frac{r}{k-r-1}}\sqrt{\sum_{j>r}\sigma_{j}({\mathbf{X}})}\leq\sqrt{2}\|{\mathbf{X}}-{\mathbf{X}}_{0}\|_{F}\leq\sqrt{2}\varepsilon\|{\mathbf{X}}\|_{F},$
using that $k\geq 2r+1$ in the second inequality. Combining this with Theorem 2, we have
${\mathbb{E}}\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|_{F}\leq{\color[rgb]{0,0,0}c_{\lambda}}(\|{\mathbf{X}}-{\mathbf{X}}_{0}\|_{F}+{\mathbb{E}}\|{\mathbf{X}}-P_{{\mathbf{A}}^{T}}{\mathbf{X}}\|_{F})\leq{\color[rgb]{0,0,0}c_{\lambda}}(1+\sqrt{2})\varepsilon\|{\mathbf{X}}\|_{F},$
where ${\color[rgb]{0,0,0}c_{\lambda}}=\max(2/\lambda,6,2\lambda+2).$
Remark 3.
It is easy to see that the oversampling condition $k>2r$ can be relaxed to any $k>r+1$ by suitably increasing the constant factor $\sqrt{2}$. Notwithstanding this factor, we see that if $\mathbf{X}$} has an exact NMF decomposition of rank $r$ and $k>r+1$ then the error of the optimal solution to the compressed problem must be also zero, comparable with the result of Theorem 1.
A high probability deviation bound for the loss $\|{\mathbf{X}}-{\mathbf{Q}}{\mathbf{Q}}^{T}{\mathbf{X}}\|_{F}$ is also known [15, Theorem 10.7]. It implies a high probability estimate for (10) in a straightforward way. Instead of Gaussian initial sketching, one can employ subsampled random Fourier transform [15] or other cost-efficient matrices [34, 11]. Moreover, similar results are available for the spectral norm instead of Frobenius (e.g., [15, Theorem 10.3]), and the results of Theorem 2 as well as Corollary 1 extend to the spectral norm loss. An interesting question for the future work is how to extend similar results to the other standard losses appearing in the applications of the nonnegative matrix factorization problem, such as $\beta$-divergence [10] or $L_{2,1}$-norm loss [22].